96. Fahrenheit temperature. Water boils at 212 degrees Fahrenheit.
a) Use the accompanying graph to determine the Celsius temperature at which water boils.
b) Find the Fahrenheit temperature of hot tap water at 70 degrees C by solving the equation.
70 = 5/9 (F – 32)

Answer by checkley75(3666)

C=5/9(212-32)
C=5/9*180
C=5*180/9
C=5*20
C=100 DEGREES IS THE ANSWER FOR BOILING WATER.
---------------------------------------------------------------
70=5/9(F-32)
70=(5F-160)/9
70*9=5F-160
5F=630+160
5F=790
F=790/5
F=158 DEGREES IS THE ANSWER.